### v1

# Increasing the clock speed of a thermodynamic computer by adding noise

Abstract
We describe a proposal for increasing the effective clock speed of a thermodynamic computer, by altering the interaction scale of the units within the computer and introducing to the computer an additional source of noise. The resulting thermodynamic computer program is equivalent to the original computer program, but runs at a higher clock speed. This approach offers a way of increasing the speed of thermodynamic computing while preserving the fidelity of computation.
I Introduction
In classical computing, the energy scales of even the smallest devices, such as transistors and gates, are large compared to that of the thermal energy, $k_{\rm B}T$. As a result, there is a clear separation of scales between signal and noise, enabling deterministic computation. This determinism comes with an energy cost: classical computers operate far above the limits of thermodynamic efficiency, and require large amounts of power and heat dissipation to ensure their reliability Bennett (1982); Landauer (1991); Ceruzzi (2003). As devices shrink to the nanoscale, the energy associated with computation becomes comparable to that of thermal fluctuations, making computation more energy efficient Frank (2017); Wolpert (2019); Conte et al. (2019). However, the comparable scales of signal and noise makes deterministic computation challenging: efficient time-dependent protocols are required to do even the simplest logic operations on the $k_{\rm B}T$-scale Gingrich et al. (2016); Dago et al. (2021); Proesmans et al. (2020a); Zulkowski and DeWeese (2014); Proesmans et al. (2020b); Blaber and Sivak (2023); Barros et al. (2024).
Fields such as thermodynamic computing view thermal fluctuations as a resource rather than a problem Conte et al. (2019); Hylton (2020). Thermodynamic computing makes use of the tendency of physical systems to evolve toward thermal equilibrium to do computation. In this framework, the thermal bath contributes to computation by providing the fluctuations necessary for state changes, and in some modes of operation the signal is the noise, with the equilibrium fluctuations of the degrees of freedom of the thermodynamic computer being the output of the calculation Aifer et al. (2023); Melanson et al. (2023).
One challenge faced by thermodynamic computing is the requirement to attain thermodynamic equilibrium. Fields such as molecular self-assembly and glass physics reveal that although we may know the thermodynamic ground state of a physical system, there is no guarantee that a system will attain equilibrium on practically-realizable timescales Hagan et al. (2011); Whitelam and Jack (2015); Biroli and Garrahan (2013); Arceri et al. (2022). The emergence of slow dynamics and kinetic traps can cause a system to spend a considerable amount of time in a nonequilibrium state. With some thermodynamic computers Aifer et al. (2023); Melanson et al. (2023) resembling spin glasses Mézard et al. (1984), which can be very slow to equilibrate, strategies to attain equilibrium as rapidly as possible are important for realizing the potential of thermodynamic computing.
In this paper we propose a way to accelerate the dynamical evolution, and hence rate of equilibration, of the class of thermodynamic computers that evolve according to overdamped Langevin dynamics in the presence of Gaussian white noise. By altering the interaction scale of the units within the thermodynamic computer, and introducing to the system an additional source of Gaussian white noise whose variance is precisely chosen, we can effectively increase the clock speed of a thermodynamic computer with no impact on its computational ability. In other words, the efficiency of noise-dominated computing can be increased by adding noise. This counter-intuitive notion emphasizes a fundamental difference between thermodynamic computing and classical forms of computing (including analog computing) whose goal is to overcome or suppress noise.
In Section II we demonstrate this clock-acceleration procedure in general, using analytic methods. In Section III we provide a simple numerical illustration of the procedure, using a thermodynamic computer program to invert a $4\times 4$ symmetric positive definite matrix. The equilibration time of even this small system can vary by orders of magnitude as the matrix is changed. This broad distribution of timescales indicates the need for a clock-acceleration procedure, because it would be inconvenient for the run time of a thermodynamic computer program to vary by orders of magnitude as its parameters are varied. We conclude in Section IV.
II Increasing the clock speed of a thermodynamic computer
We consider a thermodynamic computer composed of $N$ units with real-valued scalar amplitudes $S_{i}$. Amplitudes could represent physical distances, if units are realized mechanically Dago et al. (2021), or voltage states, if realized by RLC circuits Aifer et al. (2023); Melanson et al. (2023). Units interact via the potential $V({\bm{S}})$, which typically consists of pairwise couplings between units Aifer et al. (2023); Melanson et al. (2023).
Units evolve according to the overdamped Langevin dynamics
$\dot{S}_{i}=-\mu\frac{\partial V({\bm{S}})}{\partial S_{i}}+\sqrt{2\mu k_{\rm B}T}\,\eta_{i}(t).$ (1)
Here $\mu$, the mobility, sets the time constant of the computer. For the thermodynamic computers described in Refs. Aifer et al. (2023); Melanson et al. (2023), which are realized by RLC circuits, $1/\mu\propto RC\sim 1$ microsecond. The combination $k_{\rm B}T$ is the thermal energy scale – $k_{\rm B}$ is Boltzmann’s constant and $T$ is temperature – and $\eta$ is a Gaussian white noise with zero mean, $\langle\eta_{i}(t)\rangle=0$, and covariance $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The Kronecker delta $\delta_{ij}$ indicates independence between different noise components, and the Dirac delta $\delta(t-t^{\prime})$ indicates the absence of time correlations. The noise $\eta$ represents the thermal fluctuations inherent in the system. We consider the interaction $V$ and the noise $\eta$ to specify the program of the thermodynamic computer.
Now say that we have a thermodynamic computer program (1) whose dynamics is slow, and so takes a long time to run (and so to equilibrate, if that is our goal). The most direct way to speed up the computer is to increase the mobility parameter $\mu$, which sets the computer’s time constant. However, for RLC circuits we do not have unlimited freedom to do this, and timescales much shorter than microseconds are difficult to arrange.
An alternative is to consider running an accelerated computer program whose solution is identical to that of the original program but which takes less time to run. We can construct such a program as follows. Rescale the interaction term $V({\bm{S}})$ by a factor $\lambda\geq 1$, and introduce to the system an additional source of Gaussian white noise. If the new noise has variance $\sigma^{2}$ then the new program is described by the equation
$\dot{S}_{i}=-\mu\frac{\partial[\lambda V({\bm{S}})]}{\partial S_{i}}+\sqrt{2\mu k_{\rm B}T}\,\eta_{i}(t)+\sigma\zeta_{i}(t),$ (2)
where $\langle\zeta_{i}(t)\rangle=0$ and $\langle\zeta_{i}(t)\zeta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The added noise $\zeta$ can be thermal (from a different heat bath) or athermal in nature, provided that it is Gaussian and has no temporal correlations. The two noise terms in (2) constitute an effective Gaussian white noise
$\sqrt{2\mu k_{\rm B}T+\sigma^{2}}\,\xi_{i}(t),$ (3)
where $\langle\xi_{i}(t)\rangle=0$ and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. If we then set
$\sigma^{2}=2\mu k_{\rm B}T(\lambda-1),$ (4)
Eq. (2) can be written
$\dot{S}_{i}=-\tilde{\mu}\frac{\partial V({\bm{S}})}{\partial S_{i}}+\sqrt{2\tilde{\mu}k_{\rm B}T}\,\xi_{i}(t),$ (5)
where $\tilde{\mu}\equiv\lambda\mu$. Eq. (5) is Eq. (1) with mobility parameter rescaled by a factor $\lambda\geq 1$ (the correlations of $\eta$ and $\xi$ are the same). Thus the accelerated computer program, (2) and (4), runs the original computer program (1) but faster  [^1] A similar effect could be achieved by rescaling $V\to\lambda V$ and increasing temperature $T\to\lambda T$, but we assume this is less practical than adding an external source of noise.. Another way to see this is to rescale time $t\to\lambda t$ in (1); the result is Eq. (5). The accelerated program runs the original program with the clock speed increased by a factor of $\lambda\geq 1$. The thermodynamic behavior of the system is unchanged.
III Numerical illustration of accelerated matrix inversion
To illustrate this clock-acceleration procedure we consider the matrix inversion program of Refs. Aifer et al. (2023); Melanson et al. (2023). Here the computer’s degrees of freedom possess the pairwise interaction
$V({\bm{S}})=\sum_{\langle ij\rangle}J_{ij}S_{i}S_{j},$ (6)
where the sum runs over all distinct pairs of interactions. We choose an interaction matrix $J_{ij}$ that is positive definite and symmetric, and so has all eigenvalues non-negative and $N(N+1)/2$ distinct entries. This interaction is shown schematically in the inset of Fig. 1(a). In thermal equilibrium the probability distribution of the computer’s degrees of freedom is $\rho_{0}({\bm{S}})={\rm e}^{-\beta V({\bm{S}})}/\int{\rm d}{\bm{S}}^{\prime}{\rm e}^{-\beta V({\bm{S}}^{\prime})}$, where $\beta^{-1}\equiv k_{\rm B}T$, and so for the choice (6) we have $\langle S_{i}\rangle_{0}=0$ and
$\langle S_{i}S_{j}\rangle_{0}=\beta^{-1}(J^{-1})_{ij}.$ (7)
Here the brackets $\langle\cdot\rangle_{0}$ denote a thermal average, and $(J^{-1})_{ij}$ denotes the elements of the inverse of the matrix $J_{ij}$. Thus we can invert the matrix $J_{ij}$ by sampling the thermal fluctuations, and specifically the two-point correlations, of the units ${\bm{S}}$ in thermal equilibrium.
Figure 1: Matrix inversion thermodynamic computer program, following Refs. Aifer et al. (2023); Melanson et al. (2023). (a) Probability distribution of the reciprocal of the smallest eigenvalue for $10^{7}$ $4\times 4$ symmetric positive definite matrices $J$. The values associated with the matrices $J_{1}$ and $J_{2}$ are marked by dots. Inset: schematic of the 4-unit thermodynamic computer used to estimate the inverses of $J_{1}$ and $J_{2}$. (b) Error $E_{1}$ [Eq. (9)] in the estimate for $J_{1}^{-1}$ using the computer program (1) run for time $t_{\rm f}$; here and subsequently we state times in units of $\mu^{-1}$. The program is run $n_{\rm s}=10^{4}$ times, from which the averages $\langle S_{i}S_{j}\rangle$ are calculated. (c) The same for the matrix $J_{2}^{-1}$. Note that the horizontal scales in (b) and (c) are different. (d) and (e): Estimates of the 10 distinct elements of $J_{1}^{-1}$ and $J_{2}^{-1}$, indexed by $k$, for the program times $t_{\rm f}$ indicated. Note that the vertical axes in (d) and (e) have different scales.
We consider the case $N=4$. In Fig. 1(a) we show the probability distribution $P$ of the reciprocal $\alpha_{\rm min}^{-1}$ of the smallest eigenvalue of $10^{7}$ symmetric positive definite $4\times 4$ matrices $J$. The matrices were generated by first creating random symmetric matrices, $B$, with 10 distinct entries, each drawn uniformly from the interval $[0,1)$. We then applied Cholesky factorization to obtain a positive-definite matrix $A=B^{\rm T}B$. To ensure that all pairwise interactions in the thermodynamic computer were of order $k_{\rm B}T$ or larger, we normalized the matrix $A$ by dividing its elements by the magnitude of its smallest entry, yielding the matrix $\beta J$. This normalization step was an arbitrary choice.
The reciprocal of the smallest eigenvalue of the matrix $J$ is a rough estimate of the relaxation time of the resulting thermodynamic computer program. We chose two matrices, $J_{1}$ and $J_{2}$, whose reciprocal smallest eigenvalues are 0.011 and 0.239 respectively, and so we expect the program specified by $J_{2}$ to take at least an order of magnitude longer than that specified by $J_{1}$. We shall see that this is the case (and also that the smallest eigenvalue is not a precise estimate of the equilibration time for two-point correlations).
The matrix $\beta J_{1}$ is
$\beta J_{1}=\begin{pmatrix}33.4484&-2.13458&-27.2213&-13.3774\\ -2.13458&91.0349&1&6.58694\\ -27.2213&1&78.3708&-12.075\\ -13.3774&6.58694&-12.075&55.602\end{pmatrix},$
and its inverse is
$\beta^{-1}J_{1}^{-1}=10^{-3}\begin{pmatrix}54.806&-0.253332&21.8054&17.9513\\ -0.253332&11.0905&-0.456606&-1.47396\\ 21.8054&-0.456606&21.8886&10.0538\\ 17.9513&-1.47396&10.0538&24.6619\end{pmatrix}.$
The matrix $\beta J_{2}$ is
$\beta J_{2}=\begin{pmatrix}1.18752&1&1.00191&1.08641\\ 1&1.102484&1.02245&1.01064\\ 1.00191&1.02245&1.06828&1.03442\\ 1.08641&1.01064&1.03442&1.07344\end{pmatrix},$
and its inverse is
$\beta^{-1}J_{2}^{-1}=\begin{pmatrix}16.9091&-1.08542&11.3872&-27.0648\\ -1.08542&8.37908&-6.37593&-0.646175\\ 11.3872&-6.37593&25.497&-30.092\\ -27.0648&-0.646175&-30.092&57.9299\end{pmatrix}.$
We begin with the units of the thermodynamic computer set to zero, ${\bm{S}}={\bm{0}}$ (in thermal equilibrium we have $\langle{\bm{S}}\rangle_{0}={\bm{0}}$). We run the program (1) for time $t_{\rm f}$, and then record the value of all distinct pairwise products $S_{i}S_{j}$. To carry out these simulations we used the first-order Euler discretization of (1),
$S_{i}(t+\Delta t)=S_{i}(t)-\mu\Delta t\frac{\partial V({\bm{S}})}{\partial S_{i}}+\sqrt{2\mu\Delta tk_{\rm B}T}\,X,$ (8)
where $\Delta t=10^{-4}\mu^{-1}$ is the time step and $X\sim{\mathcal{N}}(0,1)$ is a Gaussian random number with zero mean and unit variance.
We then reset the units of the computer to zero, and repeat the process, obtaining $n_{\rm s}=10^{4}$ sets of measurements in all. On a single device an alternative would be to run one long trajectory and sample periodically (note that the resetting procedure could be run in parallel, if many devices are available). We compute the average $\langle S_{i}S_{j}\rangle$ of all distinct pairwise products over all samples. If $t_{\rm f}$ is long enough for the computer to come to equilibrium, then $\langle S_{i}S_{j}\rangle\approx\langle S_{i}S_{j}\rangle_{0}$ (for sufficiently large $n_{\rm s}$), and these measurements will allow us to recover the inverse of the relevant matrix via Eq. (7).
As a measure of the error $E$ in the thermodynamic computer’s estimate of the elements $\beta^{-1}(J^{-1})_{ij}$, we use the Frobenius norm
$E=\sqrt{\sum_{i,j}\left[\langle S_{i}S_{j}\rangle-\beta^{-1}(J^{-1})_{ij}\right]^{2}}.$ (9)
In Fig. 1(b) we show the error $E_{1}$ in the estimate of the inverse $J_{1}^{-1}$, upon using the matrix $J_{1}$ in (6) and running the program (1) for time $t_{\rm f}$ (in units of $\mu^{-1}$; subsequently, we will omit the units $\mu^{-1}$ when discussing timescales). For each run time, $n_{\rm s}=10^{4}$ samples are collected. The plot indicates that thermodynamic equilibrium is attained on timescales $t_{\rm f}\gtrsim 0.2$.
In Fig. 1(c) we show the corresponding plot for the matrix $J_{2}$. In this case, equilibrium is attained for run times $t_{\rm f}\gtrsim 200$ meaning that the program times for inverting the matrices $J_{1}$ and $J_{2}$ differ by about three orders of magnitude. The origin of this difference can be understood from the relative scales of the matrix elements involved: $J_{1}$ encodes a thermodynamic driving force of order 10s of $k_{\rm B}T$ (leading to rapid changes of ${\bm{S}}$), and the associated equilibrium fluctuations of the units ${\bm{S}}$ are on the scale of $10^{-2}$ (which take little time to establish). By contrast, $J_{2}$ encodes a thermodynamic driving force of order $k_{\rm B}T$ (leading to slower changes of ${\bm{S}}$), and some of the associated equilibrium fluctuations are 10s of $k_{\rm B}T$ in magnitude (which take more time to establish).
In Fig. 1(d) and (e) we show the thermodynamic computer’s estimate of the 10 distinct matrix elements (indexed by the variable $k$) of $J_{1}^{-1}$ (panel (d)) and $J_{2}^{-1}$ (panel (e)), for the indicated times, upon collecting $n_{\rm s}=10^{4}$ samples. An accurate estimate of $J_{2}^{-1}$ requires times in excess of 200, or of order 0.1 milliseconds for the clock speed of the thermodynamic computer of Refs. Aifer et al. (2023); Melanson et al. (2023). Given that $10^{4}$ samples are used, the total compute time would be of order a second, for a single $4\times 4$ matrix (scaling estimates Aifer et al. (2023); Melanson et al. (2023) suggest that the characteristic time for matrix inversion using an overdamped thermodynamic computer will grow as the square $N^{2}$ of the matrix dimension $N$, but even matrices of equal size can have substantially different equilibration times).
Figure 2: Accelerated matrix inversion program. (a) Error $E_{2}$ [Eq. (9)] in the estimate of the matrix $J_{2}^{-1}$ using the accelerated computer program (2) and (4), run for time $t_{\rm f}=1$ as a function of the clock-acceleration parameter $\lambda$ (black dashed line). This result is overlaid on the data of Fig. 1(c), derived from the original program (1) run for time $t_{\rm f}$ (green). We collect $n_{\rm s}=10^{4}$ samples for each program. As expected from the comparison of (1) and (5), the accelerated program run for $t_{\rm f}$ is equivalent to the original program run for $\lambda t_{\rm f}$. (b) Exact elements of $J_{2}^{-1}$ and those estimated using the accelerated program run for time $t_{\rm f}$ at two values of $\lambda$. The case $\lambda=1$ is equivalent to the original program.
To speed up the computation of $J_{2}^{-1}$ we run the accelerated program specified by Eqns. (2) and (4). The matrix $J_{2}$ is therefore replaced by the matrix $\lambda J_{2}$, and we introduce an additional source of noise with variance (4). In Fig. 2(a) we show the error $E_{2}$ in the computer’s estimate of the elements of $J_{2}^{-1}$, as a function of the clock acceleration parameter $\lambda$. Here the program time is set to $t_{\rm f}=1$, and $n_{\rm s}=10^{4}$ samples are collected. We overlay this result on that of Fig. 1(c), the error associated with the original program (1) run for time $t_{\rm f}$. As expected from comparison of (1) and (5), these results are similar: the accelerated program run for $t_{\rm f}$ is equivalent to the original program run for $\lambda t_{\rm f}$.
Using the accelerated program with $\lambda=1000$, the total compute time for inverting $J_{2}$ at the clock speeds of the computer of Refs. Aifer et al. (2023); Melanson et al. (2023) would be reduced from of order 1 second (for the original program) to of order 1 millisecond. If all samples were run on parallel hardware, the wall time for the computation would be of order 0.1 microseconds.
In panel (b) we show the inverse matrix elements extracted from the accelerated program (2) and (4) for two values of $\lambda$. As expected from the comparison of (1) and (5), the elements extracted using $\lambda=1000$ for program time $t_{\rm f}=1$ are as accurate as those extracted using the original program run for time $t_{\rm f}=1000$.
IV Conclusions
Nanoscale computer components must operate in the presence of thermal fluctuations. The premise of thermodynamic computing is to use these fluctuations, as a means of state change and as a form of signal. Here we have proposed a way of accelerating the rate of such computation by adding noise to the computer and scaling its inter-component connections appropriately.
We have shown that a thermodynamic computer realized by overdamped Langevin dynamics in the presence of Gaussian white noise, described by Eq. (1), can be accelerated by rescaling the interactions between its components and adding noise, according to Eqns. (2) and (4). The resulting dynamics, described by Eq. (5), is equivalent to the original dynamics with the basic rate of computation or clock speed increased by a factor $\lambda\geq 1$. This acceleration of the clock leaves the thermodynamic properties of the computer unchanged.
How large $\lambda$ can be made will depend on the details of the thermodynamic computer and the nature of the injected noise source. This could be thermal or athermal; it could involve injecting voltage noise using an external circuit, or digitally simulated and fed into the circuit through a programmable signal generator, or perhaps by using ‘waste noise’ from other sources. The injected noise and the rescaling of the computer’s interactions must be balanced according to (4). We have illustrated this clock acceleration numerically, using the matrix inversion program of Refs. Aifer et al. (2023); Melanson et al. (2023).
In this paper we have considered overdamped dynamics, but thermodynamic computers can also be realized by underdamped dynamics Aifer et al. (2023); Melanson et al. (2023), described generically by
$m\ddot{S}_{i}+\gamma\dot{S}_{i}=-\frac{\partial V({\bm{S}})}{\partial S_{i}}+\sqrt{2\gamma k_{\rm B}T}\,\eta_{i}(t).$ (10)
The first term represents inertial effects ($m$ is mass), the second term is the damping force ($\gamma=1/\mu$ is the damping coefficient), and the noise term represents thermal fluctuations inherent in the system. Rescaling time by a factor $\lambda\geq 1$ in (10) results in the equation
$m\ddot{S}_{i}+\lambda\gamma\dot{S}_{i}=-\lambda^{2}\frac{\partial V({\bm{S}})}{\partial S_{i}}+\sqrt{2\lambda^{3}\gamma k_{\rm B}T}\,\eta_{i}(t).$ (11)
We can produce (11) from (10) – and so effectively increase the clock speed of the computer – by adding to (10) a Gaussian white noise with zero mean and variance $2(\lambda^{3}-1)\gamma k_{\rm B}T$, rescaling $V$ by $\lambda^{2}$, and adding an additional damping term $(1-\lambda)\gamma\dot{S}_{i}$ to the right-hand side of (10). The latter requirement makes constructing an accelerated program for an underdamped thermodynamic computer more involved than doing so for an overdamped computer, but it is possible in principle.
Acknowledgments — I thank David Sivak and Corneel Casert for discussions. This work was performed at the Molecular Foundry at Lawrence Berkeley National Laboratory, supported by the Office of Basic Energy Sciences of the U.S. Department of Energy under Contract No. DE-AC02–05CH11231.